OpenStudy (joshoyen):

use the information provided to write the vertex form equation of each parabola. 1. y = -7x^2 - 14x - 6

OpenStudy (solomonzelman):

factor the first two terms out of -7, please.

OpenStudy (solomonzelman):

Note that: \(\color{#000000 }{ \displaystyle -7x^2=(-7)\cdot x^2 }\) and, \(\color{#000000 }{ \displaystyle -14x=(-7)\cdot 2x }\)

OpenStudy (solomonzelman):

can you factor \(\color{#000000 }{ \displaystyle -7x^2-14x }\) out of -7?

OpenStudy (joshoyen):

trying

OpenStudy (solomonzelman):

sure ...

OpenStudy (joshoyen):

@SolomonZelman hm, im confused

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle ab+ac=a\cdot (b+c) }\) did you know?

OpenStudy (joshoyen):

Yeah I've seen the formula before but never really knew how to use it

OpenStudy (solomonzelman):

Well, think about 2•4 + 8•4 = 4•(8+2) (factored out of 4) and in fact if you work both sides, 8 + 32 = 4 • (10) 40 = 40

OpenStudy (solomonzelman):

So if you see two components that are both products, and both have some common term "a", then this "a" can be factored out like I did with "a" in the formla, and with "4" in the example.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle -7x^2 - 14x}\) So wouldn't it make sense to say; \(\color{#000000 }{ \displaystyle (-7)\cdot x^2 + (-7)\cdot 2x}\) \(\color{#000000 }{ \displaystyle (-7)(x^2 +2x)}\)

OpenStudy (joshoyen):

ohhh okay i see now

OpenStudy (solomonzelman):

So, \(\color{#000000 }{ \displaystyle y = -7x^2 - 14x - 6}\) \(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x)-6}\)

OpenStudy (solomonzelman):

Ok, and do you know that: \(\color{#000000 }{ \displaystyle (x+a)^2=x^2+2ax+a^2}\) ?

OpenStudy (joshoyen):

hmm, no haven't seen that formula before

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle (x+a)^2=}\) \(\color{#000000 }{ \displaystyle (x+a) \cdot (x+a)=}\) \(\color{#000000 }{ \displaystyle x\cdot (x+a)+a\cdot (x+a)=}\) \(\color{#000000 }{ \displaystyle x\cdot x+x\cdot a+a\cdot x+a\cdot a=}\) \(\color{#000000 }{ \displaystyle x^2+xa+ax+a^2=}\) (xa is same as ax) \(\color{#000000 }{ \displaystyle x^2+2ax+a^2.}\)

OpenStudy (solomonzelman):

I am doing a quick derivation of that...

OpenStudy (solomonzelman):

you can look more into detail about \(\color{#000000 }{ \displaystyle x^2+2ax+a^2=(x+a)^2}\) but for know, you need to get \(x^2+2x\) into that form.

OpenStudy (solomonzelman):

you can see that in our case "a" is? (Hint: 2ax in the formula coressponds to 2x in our case)

OpenStudy (joshoyen):

a is -7

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x^2+2ax+a^2=(x+a)^2}\) \(x^2+2x\) 2x=2ax a=1

OpenStudy (solomonzelman):

1²=1

OpenStudy (solomonzelman):

So you need to add 1 to x²+2x to make it into the form of \(\color{#000000 }{ \displaystyle x^2+2ax+a^2~~~~~~~~~~~=(x+a)^2}\)

OpenStudy (solomonzelman):

but we can't just add 1, because that changes the value however we can; \(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x)-6}\) \(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x+1-1)-6}\) The rule is: \(\color{#000000 }{ \displaystyle (A)(C-D)=AC-AD}\) (you can show this, you play with 5(5-3) = 5•5 - 5•3)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x+1-1)-6}\) in our case \(\color{#000000 }{ \displaystyle x^2 +2x+1}\) is C and -1 is -D. (or [the last] 1 is D)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x+1)-(-7)(1)-6}\) \(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x+1)+7-6}\) \(\color{#000000 }{ \displaystyle y=(-7)(x^2 +2x+1)+1}\)

OpenStudy (solomonzelman):

notice that (x+1)²=x^2+2(1)(x)+1² = x² + 2x +1 so re-write the part in parenthesis.

OpenStudy (solomonzelman):

and then the vertex of \(\color{#000000 }{ \displaystyle y=a(x-h)^2+k}\) is (h,k) and equivalent the vertex of \(\color{#000000 }{ \displaystyle y=a(x+t)^2+k}\) is (-t,k)

OpenStudy (joshoyen):

oh man thank you dude, i'm gonna have to read through all this again, i just can't get it

OpenStudy (solomonzelman):

yes, and ask other people questions about anything you don't understand ... to ask is most important.

OpenStudy (solomonzelman):

Good luck

OpenStudy (whpalmer4):

Here's a slightly different approach: Your equation is a parabola: \[y = -7x^2 - 14x - 6\] If we have a parabola written in the form \[y = ax^2 + bx + c\]where \((a,b,c)\) are all constants as they are here (\(a=-7,\ b=-14,\ c=-6)\) we can find the vertex easily by knowing that the \(x\)-value of the vertex (which we will call \(h\)) will be\[x = -\frac{b}{2a}\]and the \(y\)-value (which we will call \(k\)) we get by just plugging the value of \(x\) into the formula of the parabola. Once you know the vertex is at \((h,k)\) you can write the formula for the parabola in vertex form using this: \[y = a(x-h)^2 + k\] The value of \(a\) is the same value we had for \(a\) when we found the vertex.